# DISC: Dynamic Decomposition Improves LLM Inference Scaling

### v1

@@ Abstract
Inference scaling methods often rely on decomposing problems into steps (or groups of tokens), followed by sampling and selecting the best next steps. However, these steps and their sizes are often predetermined or manually designed based on domain knowledge. We propose dynamic decomposition, a method that adaptively and automatically fractions solution and reasoning traces into manageable steps during inference. By more effectively allocating compute – particularly through subdividing challenging steps and prioritizing their sampling – dynamic decomposition significantly improves inference efficiency. Experiments on benchmarks such as APPS, MATH, and LiveCodeBench demonstrate that dynamic decomposition outperforms static approaches, including token-level, sentence-level, and single-step decompositions. These findings highlight the potential of dynamic decomposition to improve a wide range of inference scaling techniques.
Machine Learning, LLMs, Inference scaling, test-time compute, planning, reasoning, decomposition
@@ 1 Introduction
Scaling inference efficiency remains a fundamental challenge for large language models (LLMs). Many existing approaches improve inference by decomposing problems into smaller steps and systematically exploring different solutions (Feng et al., 2023; Zeng et al., 2024; Wu et al., 2024; Nori et al., 2024; Snell et al., 2024; Brown et al., 2024; Gandhi et al., 2024; Lee et al., 2025; Light et al., 2024a; Wang et al., 2025).
Some decomposition methods rely on domain-specific heuristics and hand-crafted rules (Yao et al., 2024; Zelikman et al., 2023; Zhou et al., 2022). However, manually partitioning problems or designing task-specific heuristics is costly and lacks generalization. Moreover, identifying critical steps for an LLM can be non-trivial for humans. As shown in Sec. 3.5, LLMs may assign importance to seemingly trivial words (e.g., therefore or which), which, while counterintuitive to humans, play a crucial role in autoregressive generation (Lin et al., 2025). Other approaches employ fixed, uniform step sizes, such as token- or sentence-level decomposition (Feng et al., 2023; Guo et al., 2025). All these methods rely on static decomposition strategies, where step sizes are predefined or determined via heuristics. Such rigidity wastes compute on steps that are easy for the LLM (but potentially difficult for humans) while undersampling more challenging steps.
To overcome these limitations, we propose DISC (Dynamic decomposition Improves Scaling Compute), a recursive inference algorithm that dynamically partitions solution steps based on difficulty. Unlike prior methods, DISC adapts decomposition granularity during inference based on both the available budget and problem complexity, ensuring finer granularity for more difficult steps. By leveraging the autoregressive nature of LLMs, DISC efficiently locates difficult steps through binary partitioning, focusing compute on challenging regions rather than wasting resources on trivial steps. DISC is generalizable and requires no human supervision, domain-specific heuristics, prompt engineering, or process annotations, making it widely applicable across tasks.
@@
  • We introduce DISC, a method for recursive partitioning and decomposing solutions during inference without human supervision, domain-specific heuristics, or process reward models.
  • We demonstrate how DISC integrates decomposition with inference-time search, allocating compute to high-impact, difficult steps.
  • We show that DISC improves inference scaling in terms of both sample efficiency and token efficiency.
  • We provide insights into how LLMs reason and plan by identifying critical steps in their generation process.
@@ 2 Preliminaries
Figure 1: Comparison of different automatic decomposition methods based on step size determination.
@@ 2.1 Problem Setting
We consider a reasoning and code generation setting where a dataset $\mathcal{X}=\{\boldsymbol{x}^{(i)}\}_{i=1}^{N}$ consists of problem prompts $\boldsymbol{x}$, and a reward model $R:\mathcal{X}\cdot\mathcal{Y}\rightarrow[0,1]$ evaluates generated solutions $\boldsymbol{y}\in\mathcal{Y}$. This includes program synthesis, where correctness is verified using ground-truth tests (Chen et al., 2021; Austin et al., 2021), and mathematical reasoning, where solutions are validated numerically (Hendrycks et al., 2021a; Cobbe et al., 2021). The reward model can be a ground-truth verifier, a trained heuristic (Zhang et al., 2024), self-consistency (Wang et al., 2023a), or an LLM-as-a-judge (Zheng et al., 2023). Since our focus is on step decomposition rather than verification, we use the ground-truth reward model where available. We assume access to a pretrained language model $\pi$ that generates text autoregressively. A generated response $\boldsymbol{y}$ consists of both the final solution and the reasoning chain leading to it, and can be represented as a sequence of tokens $\boldsymbol{y}=(y_{0},...,y_{L_{\boldsymbol{y}}})$. Additionally, solutions can be partitioned into solution steps $\boldsymbol{y}=(\boldsymbol{y}_{0},...,\boldsymbol{y}_{K})$, where each step $\boldsymbol{y}_{i}$ is a contiguous string of tokens. A partial solution up to step $k$ is defined as $\boldsymbol{y}_{1...k}:=\boldsymbol{y}_{1}\cdot\boldsymbol{y}_{2}\cdot...\cdot\boldsymbol{y}_{k}$, and its rollout or completion, denoted $\boldsymbol{y}_{1...k+}$, is the continuation generated by $\pi$ until an end-of-sequence token (EOS). The size of a solution step, $|\boldsymbol{y}_{i}|$, refers to its length in tokens or characters.
2.2 Prior Automatic Decomposition Methods
@@
Single-step generation. In a single-step generation, the entire solution is generated in one pass from the prompt to the EOS token, treating it as a single action. This approach underlies the widely used inference scaling method best of n (BoN) (Cobbe et al., 2021; Lightman et al., 2023; Snell et al., 2024; Liang et al., 2024), where $n$ complete solutions are sampled, and the highest-scoring one is selected. Single-step generation also plays a role in alignment and fine-tuning methods such as DPO (Rafailov et al., 2024) and RLOO (Ahmadian et al., 2024).
@@
Token-level decomposition. At the opposite end of the spectrum, token-level decomposition treats each atomic token as an individual step. While this approach dramatically increases search complexity, it enables fine-grained search that can yield higher performance gains given sufficient compute (Feng et al., 2023).
@@
Newline and sentence-level decomposition. A commonly used decomposition method segments LLM generations into sentences or lines based on delimiters such as periods or newlines (Hao et al., 2023; Feng et al., 2023; Yao et al., 2024). Typically, each newline corresponds to a new paragraph, equation, or line of code, which often encapsulates a distinct reasoning step.
Problem: Automatic and scalable decomposition Existing decomposition methods are task-specific, manual, and static, limiting their adaptability and scalability.
2.3 Step Sampling
Inference-time scaling methods must balance exploration at the current step with exploration of future steps. We implement a simple dynamic sampling process, referred to as negative binomial sampling, where we continue sampling completions until the sum of their rewards exceeds a predefined threshold $\sigma$. More formally, the number of samples $M$ drawn from a partial solution $\boldsymbol{y}_{1...k}$ is the smallest integer satisfying $\sum_{i=1}^{M}R(\boldsymbol{x}\cdot\boldsymbol{y}_{1...k+}^{(i)})\geq\sigma$.
Where $\boldsymbol{y}_{1...k+}^{(i)}$ represents the $i$-th sampled completion from the partial solution. This process ensures efficient allocation of compute by dynamically adjusting the number of samples per step. It achieves this by either: (a) continuing to sample until a sufficiently high-reward completion is found, or (b) stopping early when additional sampling is unlikely to yield significant improvements, thus redirecting the compute to future steps. The stopping criterion is governed by $\sigma$: when accumulated reward from completions surpasses $\sigma$, the method assumes further sampling is unnecessary. For a fair comparison, we apply this sampling method uniformly across all decomposition methods in our experiments.
2.4 Inference Scaling Methods and Decomposition
Since our study focuses on decomposition rather than search, we primarily use greedy step search as the search method. In greedy step search, multiple candidate steps are sampled at each iteration, but only the highest-scoring step is retained, while the rest are discarded. The process then repeats, conditioning future steps on the best step found so far. We also perform ablation studies comparing Monte Carlo Tree Search (MCTS)  (Feng et al., 2023; Light et al., 2024b) and beam search  (Xie et al., 2024), two commonly used inference scaling methods. These comparisons, presented in Sec. 4.2, highlight how different search strategies interact with decomposition. Additional details on MCTS and beam search are provided in App. E.

### v3

@@ Abstract
Inference scaling methods for LLMs often rely on decomposing problems into steps (or groups of tokens), followed by sampling and selecting the best next steps. However, these steps and their sizes are often predetermined or manually designed based on domain knowledge. We propose dynamic decomposition, a method that adaptively and automatically partitions solution and reasoning traces into manageable steps during inference. By more effectively allocating compute – particularly through subdividing challenging steps and prioritizing their sampling – dynamic decomposition significantly improves inference efficiency. Experiments on benchmarks such as APPS, MATH, and LiveCodeBench demonstrate that dynamic decomposition outperforms static approaches, including token-level, sentence-level, and single-step decompositions, reducing the pass@10 error rate by 5.0%, 6.7%, and 10.5% respectively. These findings highlight the potential of dynamic decomposition to improve a wide range of inference scaling techniques.
@@ 1 Introduction
Scaling inference efficiency remains a fundamental challenge for large language models (LLMs). Many existing approaches improve inference by decomposing problems into smaller steps and systematically exploring different solutions (feng2023alphazero, ; zeng2024scaling, ; wu2024inferencescalinglawsempirical, ; nori2024medprompto1explorationruntime, ; snell2024scaling, ; brown2024large, ; gandhi2024stream, ; lee2025evolvingdeeperllmthinking, ; light2024strategist, ; anonymous2025planning, ). Some decomposition methods often rely on domain-specific heuristics and hand-crafted rules (yao2024tree, ; zelikman2023parsel, ; zhou2022least, ). However, manually partitioning problems or designing task-specific heuristics is costly and lacks generalization. Moreover, identifying critical steps for an LLM can be non-trivial for humans. LLMs may assign importance to seemingly trivial words (e.g., therefore or which), which, while counterintuitive to humans, play a crucial role in autoregressive generation (lin2025criticaltokens, ). Other approaches employ fixed, uniform step sizes, such as token- or sentence-level decomposition (feng2023alphazero, ; guo2025deepseek, ). All these methods rely on static decomposition strategies, where step sizes are predefined or determined via heuristics. Such rigidity risks overusing compute on steps that are easy for the LLM (but potentially difficult for humans) while undersampling more challenging steps.
To overcome these limitations, we propose DISC (Dynamic decomposition Improves Scaling Compute), a recursive inference algorithm that dynamically partitions solution steps based on difficulty. Unlike prior methods, DISC adapts decomposition granularity during inference based on both the available budget and problem complexity, ensuring finer granularity for more difficult steps. By leveraging the autoregressive nature of LLMs, DISC efficiently locates difficult steps through dynamically proposing step sizes, focusing compute on challenging regions rather than wasting resources on trivial steps. DISC is generalizable, requires no human supervision, domain-specific heuristics, prompt engineering, or process annotations, and is easily parallelizable, making it widely applicable across tasks. Furthermore, DISC is plug-and-play with off-the-shelf search algorithms and can be naturally integrated with greedy, beam, or Monte Carlo Tree Search.
@@
  • We introduce DISC, a method for dynamically adjusting step sizes and decomposing solutions during inference without human supervision, domain-specific heuristics, or process reward models.
  • We demonstrate how DISC integrates decomposition directly into inference-time search, allocating compute more effectively toward high-potential solution prefixes.
  • We show that DISC improves inference scaling in terms of both sample efficiency, token efficiency, and runtime, achieving up to 10% reduction in error relative to the baselines and up to 4x increase in accuracy over the base model, with just 10 samples, including with reasoning models.
  • We provide both empirical and theoretical insights into LLM reasoning, including identifying critical intermediate steps and analysis of how DISC helps discover optimal solutions.
@@ 2 Preliminaries
Figure 1: Comparison of automatic decomposition strategies based on step size. Coarser steps accelerate the search process but risk skipping over optimal solutions and committing to suboptimal prefixes. In contrast, finer steps ensure more precise decisions but lead to slower search. A dynamic strategy that adapts step size based on LLM feedback offers a balanced approach, combining the efficiency of coarse steps with the precision of fine-grained decomposition.
@@ 2.1 Problem Setting
We consider a reasoning and code generation setting where we are given: a dataset $\mathcal{X}=\{\boldsymbol{x}^{(i)}\}_{i=1}^{N}$, a pretrained, autoregressive LLM $\pi$ that generates solutions $\boldsymbol{y}\in\mathcal{Y}$, and a reward model $R:\mathcal{X}\cdot\mathcal{Y}\rightarrow[0,1]$ that evaluates the generated solutions. The goal is to find inputs to the LLM that produce solutions with high reward. This setting includes program synthesis, where correctness is verified using ground-truth tests (chen2021evaluating, ; austin2021program, ), and mathematical reasoning, where solutions are validated numerically (hendrycks2021measuring, ; cobbe2021training, ). The reward model can be a ground-truth verifier, a trained heuristic (zhang2024generativeverifiersrewardmodeling, ), self-consistency (wang2023selfconsistency, ), or an LLM-as-a-judge (zheng2023judgingllmasajudgemtbenchchatbot, ) and because our focus is on decomposition rather than verification, we use the ground-truth reward model where available.
We use the following hierarchical token notation: $y$ is a token, $\boldsymbol{p}$ is a prefix that starts at the prompt and concatenates multiple steps, $\boldsymbol{s}$ is a suffix that concatenates multiple steps and ends with the EOS token, and $\boldsymbol{y}$ is a complete solution that starts at the prompt, concatenates multiple steps, and ends at the end of sequence token EOS. We denote a concatenation of two tokens or token sequences with $\cdot$, e.g. $\boldsymbol{p}\cdot\boldsymbol{s}$ is the concatenation of prefix $\boldsymbol{p}$ and suffix $\boldsymbol{s}$ to form a complete solution $\boldsymbol{y}$. We denote the sampled suffix from prefix $\boldsymbol{p}$ using the LLM policy as: $\boldsymbol{s}\sim\pi(\cdot|\boldsymbol{p})$. We denote an optimal solution with $\boldsymbol{y}^{*}$ and an optimal suffix as $\boldsymbol{s}^{*}$, where there exist multiple optimal solutions. For our analysis, we use the convention that $\pi(\boldsymbol{s}^{*}|\boldsymbol{p})=0$ if there does not exist a completion $\boldsymbol{s}^{*}$ such that $\boldsymbol{p}\cdot\boldsymbol{s}^{*}$ is optimal. The size of a string $|\boldsymbol{y}|$, refers to its length in tokens or characters.
2.2 Existing Decomposition Methods
@@
Single-step generation. In a single-step generation, the entire solution is generated in one pass from the prompt to the EOS token, treating it as a single action. This approach underlies the widely used inference scaling method best of n (BoN) (cobbe2021training, ; lightman2023let, ; snell2024scaling, ; liang2024improving, ), where $n$ complete solutions are sampled, and the highest-scoring one is selected. Single-step generation also plays a role in alignment and fine-tuning methods such as DPO (rafailov2024direct, ) and RLOO (ahmadian2024back, ).
@@
Token-level decomposition. At the opposite end of the spectrum, token-level decomposition treats each atomic token as an individual step. While this approach dramatically increases search complexity, it enables fine-grained search that can yield higher performance gains given sufficient compute (feng2023alphazero, ).
@@
Newline and sentence-level decomposition. A commonly used decomposition method segments LLM generations into sentences or lines based on delimiters such as periods or newlines (hao2023reasoning, ; feng2023alphazero, ; yao2024tree, ). Typically, each newline corresponds to a new paragraph, equation, or line of code, which often encapsulates a distinct reasoning step.
Challenge: Automatic and scalable decomposition Existing decomposition methods are static and manually designed, resulting in either slow convergence to good performance or fast convergence to poor performance. We propose adaptive decomposition for fast convergence to a good performance.
2.3 Search for Inference Scaling
We differentiate between decomposition and search, where decomposition controls the number of steps between new branches and search is the process of selecting which branch to explore. In other words, decomposition is the construction of nodes and edges, and search is a process that occurs on that structure. In this work, we propose a decomposition method that is plug-and-play with different search methods. In our experimental results in Sec. 4.1, we compare decomposition methods: token-level decomposition, sentence-level decomposition, and DISC with the search methods: greedy, beam (xie2024self, ), and Monte Carlo Tree Search (feng2023alphazero, ; light2024scattered, ). Implementation details are provided in App. F.
Figure 2: Multiple iterations of Alg. 1. DISC dynamically refines its step sizes across iterations, advancing and contracting the prefix at which it samples from.